sum of of age kim and pam is 20. and pam is 4 yrs younger than kim what are the ages.

Answer by nerdybill(7384)

sum of of age kim and pam is 20. and pam is 4 yrs younger than kim what are the ages.
.
Let k = age of Kim
then
k-4 = age of Pam
.
k + k-4 = 20
2k-4 = 20
2k = 24
k = 12 years (Kim)
.
Pam:
k-4 = 12-4 = 8 years